Two angles are complementary. Their difference is 45 degrees. Find the angles?
Answer by nerdybill(7384)

Two angles are complementary. Their difference is 45 degrees. Find the angles?
.
Let x = one of the two angles
then because they are complementary:
90-x = second of the two angles
.
(90-x)-x = 45
90-2x = 45
-2x = -45
x = 22.5 degrees
.
second angle:
90-x = 90-22.5 = 67.5 degrees
.
answer: 67.5 and 22.5 degrees